Shift report 20030309 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Mar 09 2003 - 23:55:37 EST

  • Next message: Ian Bearden: "Data reduction, calibrations"
    ------------------------------------------------------------
    Date: Sun Mar  9 23:55:37 2003, Shift: 16:00-24:00
    Supervisor: FR, with: SL
    Log book#14 pages 53 - 53
    
    Shift Summary:
    ------------------------------------------------------------
    
    1) Beam from 18:30 to 23:45. Rates at the beginning of
       the store: ZDC~5.6k and TOFW~14.
    
    2) Calibration runs: 8282, 8286 and 8287.
    3 data runs recorded, all for the same setting: 
    MRS  @90(A1050) and FS@4(A1/1).
      Run 8288 has 15k trig 2 events and 242k trig 3 events.
      Run 8289 has 9k trig 2 events and 134k trig 3 events.
      Run 8290 has 3.5k trig 2 events and 52k trig 3 events.   
    
    3) No detector problems noticed.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sun Mar 09 2003 - 23:56:25 EST